OpenStudy (anonymous):

Look at the formula for finding the volume of a rectangular prism below. V = lwh Which of the following is the formula to find width (w)?

OpenStudy (anonymous):

w=Vl w=v/lh w=ih/v w=lwh

OpenStudy (anonymous):

w= V/lh

OpenStudy (anonymous):

thx for the medal:)